OpenStudy (rockets31):

how would you find the slope of a line that is perpendicular to y=-3x+1; that passes through 2,2

OpenStudy (anonymous):

Perpendicular lines have slopes that multiply to -1. The line y = -3x+1 has a slope of -3. So, a line perpendicular to it will have a slope of?

OpenStudy (rockets31):

oh ok 3 then thanks had a brain fart

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

1/3 * -3 = -1

OpenStudy (rockets31):

oh ok

OpenStudy (anonymous):

Let slopes of the lines be m1 and m2. When they are perpendicular, m1 * m2 = -1. In our case, you are given m1 = -3. Since -3 * m2 = -1, m2 = 1/3.